OpenStudy (anonymous):

Trig help please? Solve the system using algebra and trigonometry. r = 2sin(theta) r = 2sin2(theta)

OpenStudy (anonymous):

sin(a+b) = sin(a)cos(b) + sin(b)cos(a) sin(2a) = sin(a+a)

OpenStudy (anonymous):

Solve the system: 2sin x = 2sin 2x -> sin x = 2*sin x*cos x = sin x*(1 - 2cos x) = 0 a) sin x = 0 = sin 0; sin Pi or sin 2Pi -> x = 0, Pi, or 2Pi b) (1 - 2cos x) = 0 -> cos x = 1/2 = cos Pi/3; and cos x = cos 5Pi/3 -> x = Pi/3 and x = 5Pi/3

OpenStudy (anonymous):

Check the system with x = Pi/3. y1 = 2sin x = 2*(V3/2) = V3 y2 = 2sin 2x = 2*sin (2PI/3) = 2(V3)/2) = V3. So, Pi/3 is a true solution.